OpenStudy (anonymous):

A rancher has 200 feet of fencing with which to enclose two adjacent rectangular corrals. What dimensions should be used so that the enclosed area will be a maximum?

OpenStudy (anonymous):



OpenStudy (anonymous):

maximation problem your given the fencing of 200 (perimeter) 2length + 2width = 200 and you want to maximize area length(width) = area solve for length length = 100 - width substitute into area (100-width)(width) = 100width - width^2 take derivative with respect to width and set = to 0 100 - 2width = 0 solve for width 100 = 2width width = 50 plug and solve for length length = 100 -50 = 50

OpenStudy (anonymous):

lol sorry i didnt see that pic let me try again

OpenStudy (anonymous):

your perimeter = 4x + 2y = 200 area = xy solve for x 4x=200-2y x = 50 - y/2 subsitute into area area = (50-y/2)(y) = 50y - y^2/2 take derivative and set = 0 50 - y = 0 y = 50 plug and solve for x x = 50 - 50/2 x = 25